# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 10.2.3 Two-dimensional multiflavor SO($N_{c}$) gauge models
We finally mention the model with $N_{f}=2$, in which the global symmetry group is the Abelian O(2) group. Two-flavor SO($N_{c}$)-gauge models behave differently, in that they undergo a finite-temperature topological BKT transition [413, 414, 415, 440, 441, 442, 443, 444, 445], between the high-temperature disordered phase and a low-temperature spin-wave phase characterized by quasi-long range order with vanishing magnetization.  [^50] We recall that BKT transitions are characterized by an exponentially divergent correlation length $\xi$ at a finite critical temperature. Indeed, $\xi$ behaves as $\xi\sim\exp(c/\sqrt{T-T_{c}})$ approaching the BKT critical temperature $T_{c}$ from the high-temperature phase. The FSS analyses of the MC results reported in Ref. [419] confirm the existence of the finite-temperature BKT transitions.
@@ 11 Effects of perturbations breaking the local gauge symmetry
Gauge symmetries can emerge as low-energy effective symmetries of many-body systems [446, 20, 447, 21]. However, because of the presence of microscopic gauge-symmetry violations, it is crucial to understand the effects of perturbations giving rise to explicit gauge-symmetry breaking (GSB). One would like to understand whether, and when, GSB perturbations destabilize the emergent gauge model— consequently, a gauge-invariant dynamics would be observed only if an appropriate tuning of the model parameters is performed—, or whether, and when, GSB perturbations do not effectively change the low-energy dynamics—in this case the gauge-symmetric theory would describe the asymptotic dynamics even in the presence of some (possibly small) violations. This issue is also crucial in the context of analog quantum simulations, for example when controllable atomic systems are engineered to effectively reproduce the dynamics of gauge-symmetric theoretical models. Indeed, in the proposed artificial gauge-symmetry realizations [448, 449, 450], the gauge symmetry in not exact and it is only expected to effectively emerge in the low-energy dynamics.
The gauge symmetry can be broken by adding two different types of operators. First, one may add a gauge-fixing term, which leaves gauge-invariant correlations invariant and does not change the physical behavior of the model, although it may allow one to observe gauge degrees of freedom that are relevant in the RG description of the critical behavior, see, e.g., Secs. 5 and 6. Here we discuss two specific GSB operators [270, 271, 451] analogous to photon-mass terms, that break the gauge invariance of LAH models with noncompact and compact U(1) gauge variables, introduced in Secs. 5 and 6 respectively. However, we expect that the main features can be straightforwardly extended to generic GSB terms.
@@ 11.1 Gauge-symmetry breaking in lattice Abelian Higgs models with noncompact gauge fields
Figure 25: Sketch of the phase diagrams for the noncompact Higgs model with GSB perturbation (118). (Top panels) Phase diagram for $\kappa=0$ (it also holds for the compact model, with $w$ replacing $m^{2}$): (a) $N=1$ (a) and (b) $N\geq 2$; scalar correlations are disordered in the D phase, tensor correlations are ordered in the T and TV phase, while vector correlations are ordered only in the TV phase. (Central panels) Phase diagram (c) for $J=\infty$ as a function of $\kappa$ and $m^{2}$, and (d) for $\kappa$ large as a function of $J$ and $m^{2}$. (Bottom panels) Phase diagrams for small $m^{2}$: (e) $N=1$, (f) $N\geq 2$. Along each transition line we report the expected corresponding universality class.
@@
$H=-2NJ\sum_{{\bm{x}},\mu}\hbox{Re}\left(e^{iA_{{\bm{x}},\mu}}\,\bar{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}+\hat{\mu}}\right)+{m^{2}\over 2}\sum_{{\bm{x}}\mu}A_{{\bm{x}},\mu}^{2}.$ (119)
For $m^{2}\to\infty$, implying $A_{{\bm{x}},\mu}=0$, one obtains a simple scalar Hamiltonian with enlarged O$(2N)$ symmetry. Indeed, if we set $z_{\bm{x}}^{a}=\sigma^{a}_{\bm{x}}+i\sigma^{a+N}_{\bm{x}}$, where $\sigma^{a}_{\bm{x}}$ is a real $2N$-component vector, we obtain
$H=-JN\sum_{{\bm{x}}\mu}\sigma_{\bm{x}}\cdot\sigma_{{\bm{x}}+\hat{\mu}}.$ (120)
@@
Taking into account the fact that for $\kappa=0$ and $m^{2}=0$ the LAH model reduces itself to a lattice CPN-1 model, see Sec. 3.2.1, and therefore there is a corresponding finite-temperature transition for $N\geq 2$, and no transition for $N=1$, one arrives at the phase diagrams reported in the two upper panels of Fig. 25. For $N=1$ [phase diagram (a)] there is a single transition line, while for $N\geq 2$ [phase diagram (b)], there are three different transition lines meeting in a single point. For small values of $m^{2}$ the GSB perturbation is not strong enough to induce long-range order in the vector correlators of ${\bm{z}}_{\bm{x}}$, while the gauge invariant spin-2 tensor operator in Eq. (57) remains polarized, as for $m^{2}=0$. We expect the CPN-1 behavior to extend along the D-T transition line, which starts at $m^{2}=0$, as shown in Ref. [451] for a general GSB perturbation of the compact model, and verified numerically in Ref. [271]. In particular, for $N=2$, continuous transitions should belong to the O(3) universality class. The irrelevance of the GSB for small $m^{2}$ is not unexpected, since gauge modes do not play any active role at the CPN-1 transition. For large $m^{2}$ values, in the low temperature phase long-range correlations emerges also in vector correlators. A natural hypothesis is that the critical behavior observed along the T-TV and D-TV transition lines, for finite values of $J$ and $m^{2}$, is the same as that observed for $J\to\infty$ and $m^{2}\to\infty$, respectively. Thus, continuous transition should belong to the XY and O($2N$) universality classes.
@@
The phase diagrams reported in the two upper panels of Fig. 25 are supposed to hold also for small values of $\kappa$. On the other hand, for large values of $\kappa$, we expect a different behavior. Before discussing the large-$\kappa$ behavior, let us consider the limit $J\to\infty$ for generic $\kappa$ and $m^{2}$. Since in this limit the gauge fields take the form (121), the effective Hamiltonian is equivalent to an IXY model (discussed in Sec. 3.5.1) coupled with the Villain XY model (122). In panel (c) of Fig. 25 we show the expected phase diagram, showing a transition line connecting the Villain XY model transition for $\kappa=0$ with the transition of the IXY model on the $m^{2}=0$ axis. Continuous transitions should belong to the XY vector universality class for $m^{2}\not=0$. While there is an XY transition if $m^{2}$ is varied for fixed small $\kappa$, no transition occurs when $\kappa$ is fixed to a large value. Thus, for large $\kappa$ we expect a single transition line, see panel (d) of Fig. 25, connecting the $m^{2}=0$ transition to the O($2N$) transition occurring for $m^{2}\to\infty$. It is natural to expect that continuous transitions along the whole line belong to the O($2N$) universality class for $m^{2}>0$. In particular, the charged transitions that occur for $m^{2}=0$ and $N>N^{*}$ should be unstable with respect to the addition of the GSB term.
@@
We finally consider the expected phase diagrams for fixed small values of $m^{2}$, see panels (e) and (f) of Fig. 25. They are qualitatively similar to the phase diagrams occurring for $m^{2}=0$, see Fig. 2. However, the nature of the phases, and therefore of the transitions, is different. The Coulomb phase is replaced by a disordered phase, in which both scalar and gauge modes are gapped, while in the low-temperature phase both scalar and gauge fields are ordered. Note that, in the latter phase, because of the absence of gauge invariance, both the gauge-invariant tensor correlations and the non-gauge-invariant vector correlations are long ranged for $N\geq 2$. In the molecular phase (which is present only for $N\geq 2$), the addition of the GSB term changes the nature of the gauge excitations—the Coulomb phase is replaced by a gapped gauge-field phase—but not that of the scalar modes: tensor correlations display long range order and vector correlations are disordered, as in the gauge-invariant model. As for the transition lines, the small-$\kappa$ transition line (which is only present for $N\geq 2$) corresponds to CPN-1 transitions as for $m^{2}=0$: the GSB term is an irrelevant perturbation. On the other hand, the nature of the other two transitions varies. The topological IXY transitions are now replaced by standard XY vector transitions, while the continuous charged transitions that occur for large values of $N$ become O($2N$) vector transitions.
@@ 11.2 Gauge-symmetry breaking in lattice Abelian Higgs models with compact gauge fields
The effects of GSB perturbations to the unit-charge ($Q=1$) LAH model with compact gauge variables, see Sec. 6.1, are discussed in Refs. [271, 451]. We consider a GSB term that is analogous to the photon-mass term in the small gauge limit, i.e.,
$H_{m}=-w\sum_{{\bm{x}}\mu}\hbox{Re }\lambda_{{\bm{x}},\mu}.$ (123)
@@
The behavior of this combined model is analogous to the one presented in the noncompact case. Indeed, for $w\to\infty$, we can set $\lambda_{{\bm{x}}\mu}=1$, and thus we obtain again the O($2N$) vector model (120). On the other hand, for $J\to\infty$, $\lambda_{{\bm{x}},\mu}=1$ modulo gauge transformations, i.e., $\lambda_{{\bm{x}},\mu}=\exp(i\phi_{\bm{x}}-i\phi_{{\bm{x}}+\hat{\mu}})$. Substituting in the GSB term (123), we obtain the effective Hamiltonian
$H_{m}=-w\sum_{{\bm{x}}\mu}\cos(\phi_{\bm{x}}-\phi_{{\bm{x}}+\hat{\mu}}),$ (125)
@@
We wish now to present an argument—it rephrases a similar argument of Ref. [41]—that shows that this result is a general property of GSB perturbations at transitions where gauge fields are not critical. Indeed, let us consider the partition function of a generic model with a GSB perturbation that only depends on the gauge fields:
$Z=\int[dU_{{\bm{x}}\mu}]\,[d\Phi_{\bm{x}}]\,\exp[-\beta H(U_{{\bm{x}},\mu},\Phi_{\bm{x}})],\qquad H(U_{{\bm{x}},\mu},\Phi_{\bm{x}})=H_{\rm GI}(U_{{\bm{x}},\mu},\Phi_{\bm{x}})+H_{\rm GSB}(U_{{\bm{x}},\mu}),$ (126)
where $H_{\rm GI}$ is gauge invariant Hamiltonian, and $H_{\rm GSB}$ is a generic Hamiltonian term breaking the gauge symmetry, depending only on the link gauge variables $U_{{\bm{x}},\mu}$. We now perform a change of variables on the scalar and gauge fields that corresponds to a gauge transformation—thus $Z$, or any expectation value of gauge-invariant operators, does not change. In particular, we redefine $U_{{\bm{x}}\mu}\to V_{{\bm{x}}}U_{{\bm{x}}\mu}V^{\dagger}_{{\bm{x}}+\mu}$. As $H_{\rm GI}$ is gauge invariant, the partition function becomes
$Z=\int[dU_{{\bm{x}}\mu}][d\Phi_{\bm{x}}]\,\exp\left[-\beta H_{\rm GI}(U_{{\bm{x}},\mu},\Phi_{\bm{x}})-\beta H_{\rm GSB}(V_{{\bm{x}}}U_{{\bm{x}},\mu}V^{\dagger}_{{\bm{x}}+\mu})\right].$ (127)
@@
The new Hamiltonian $\widehat{H}$ is clearly gauge invariant, and equivalent to the original one, if we consider the partition function and, more generally, any gauge-invariant correlation function. The Hamiltonian $\widehat{H}$ contains interactions between the gauge variables $U_{{\bm{x}},\mu}$ and $U_{{\bm{y}},\nu}$ at any distance $|{\bm{x}}-{\bm{y}}|$. However, if these interactions are exponentially suppressed for $|{\bm{x}}-{\bm{y}}|\to\infty$, then $\widehat{H}$ represents a gauge-invariant model with effective short-range interactions. This is the general scenario that we expect to emerge when the gauge fields are not critical. This can be verified by performing a strong-coupling expansion assuming $\beta w$ small. For example, in the case of the GSB perturbation (123), $\widehat{H}$ can be written as a sum of lattice loops. In the strong-coupling expansion, a lattice loop of length $L$ is weighted by a factor that behaves as $(\beta w)^{L}$ for $\beta w\to 0$. For instance, the leading term is the plaquette, with a weight of order $(\beta w)^{4}$, which renormalizes the value of $\gamma$. The next term corresponds to the $2\times 1$ plaquette, with a coefficient proportional to $(\beta w)^{6}$, and so on. Couplings therefore scale as $\exp(-a|{\bm{x}}-{\bm{y}}|)$, with $a\sim-\ln(\beta w)$, showing the short-range nature of the interactions.
This argument proves that, for small values of GSB parameter $w$, the partition function and gauge-invariant correlations can be computed in an equivalent gauge-invariant theory, without GSB terms, with short-range interactions. Finally, to conclude the argument, let us note that we are considering a model in which gauge fields do not play any role, i.e., the critical behavior is independent of the gauge-field interactions, which is the same as in the original model with $\gamma=0$. Therefore we conclude that the phase structure is independent of $w$ for sufficiently small values. Note that the argument does not rely on the Abelian nature of the theory, and thus is should also hold in non-Abelian gauge models.
It is interesting to note that the same discussion also applies to a different model without gauge fields. Indeed, for $\kappa=0$, one can integrate out the link variables $\lambda$, obtaining an effective Hamiltonian
@@
with $\widehat{w}=w/(2JN)$. As it occurs for the CPN-1 model, we expect the Hamiltonian $H_{{\rm eff}}$ to have the same critical behavior as the Hamiltonian obtained by replacing the Bessel function $I_{0}(x)$ with its small-$x$ expansion. Since

### v3

@@ 10.2.3 Lattice SO($N_{c}$) gauge models
We finally consider the model with $N_{f}=2$, in which the global symmetry group is the Abelian O(2) group. Two-flavor SO($N_{c}$)-gauge models behave differently, in that they undergo a finite-temperature topological BKT transition [457, 458, 459, 484, 485, 486, 487, 488, 489], between the high-temperature disordered phase and a low-temperature spin-wave phase characterized by quasi-long range order with vanishing magnetization. [^52] We recall that BKT transitions are characterized by a finite critical temperature with an exponentially divergent correlation length $\xi$: $\xi$ behaves as $\xi\sim\exp(c/\sqrt{T-T_{c}})$ approaching the BKT critical temperature $T_{c}$ from the high-temperature phase. The FSS analyses of the MC results reported in Ref. [463] confirm the existence of finite-temperature BKT transitions.
@@ 11 Effects of perturbations breaking the local gauge symmetry
Gauge symmetries can emerge as low-energy effective symmetries of many-body systems [490, 20, 491, 21]. However, because of the presence of microscopic gauge-symmetry violations, it is crucial to understand the effects of perturbations that explicitly break gauge invariance. One would like to understand whether, and when, gauge-symmetry breaking (GSB) perturbations destabilize the emergent gauge model— consequently, a gauge-invariant dynamics would be observed only if an appropriate tuning of the model parameters is performed—or whether, and when, GSB perturbations do not effectively change the low-energy dynamics—in this case the gauge-symmetric theory would describe the asymptotic dynamics even in the presence of some (possibly small) violations. This issue is also crucial in the context of analog quantum simulations, for example when controllable atomic systems are engineered to effectively reproduce the dynamics of gauge-symmetric theoretical models. Indeed, in the proposed artificial gauge-symmetry realizations [492, 493, 494], the gauge symmetry in not exact and it is only expected to effectively emerge in the low-energy dynamics, see, e.g., Refs [495, 496, 497, 498, 499, 500] for some experimental results.
In this section we discuss GSB effects in LAH models with noncompact and compact gauge fields. The gauge symmetry can be broken by adding different nongauge invariant terms to the gauge-invariant LAH Hamiltonian. A first peculiar class of GSB terms is that of gauge fixings: they leave gauge-invariant correlations invariant, without changing the physical gauge-invariant behavior of the model, although they may allow one to observe gauge degrees of freedom that are relevant in the RG description of the critical behavior, see, e.g., Secs. 5 and 6. Here we discuss GSB perturbations that also affect gauge-invariant correlations. Specifically, we consider operators analogous to photon-mass terms, which break the gauge invariance of LAH models introduced in Secs. 5 and 6, respectively. However, we expect the main features of the results reviewed here to be valid also in the presence of more general GSB terms.
@@ 11.1 Gauge-symmetry breaking in lattice Abelian Higgs models with noncompact gauge fields
Figure 25: Sketch of the phase diagrams for the noncompact Higgs model with GSB perturbation (121). (Top panels) Phase diagram for $\kappa=0$ (it also holds for the compact model, with $w$ replacing $m^{2}$): (a) $N=1$ and (b) $N\geq 2$; scalar correlations are disordered in the D phase, tensor correlations are ordered in the T and TV phase, while vector correlations are ordered only in the TV phase. (Central panels) Phase diagram (c) for $J=\infty$ as a function of $\kappa$ and $m^{2}$, and (d) for $\kappa$ large as a function of $J$ and $m^{2}$. (Bottom panels) Phase diagrams for small $m^{2}$: (e) $N=1$, (f) $N\geq 2$. Along each transition line we report the expected corresponding universality class.
@@
$H=-2NJ\sum_{{\bm{x}},\mu}\hbox{Re}\left(e^{iA_{{\bm{x}},\mu}}\,\bar{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}+\hat{\mu}}\right)+{m^{2}\over 2}\sum_{{\bm{x}},\mu}A_{{\bm{x}},\mu}^{2}.$ (122)
For $m^{2}\to\infty$, one has $A_{{\bm{x}},\mu}=0$, so one obtains a simple scalar Hamiltonian with enlarged O$(2N)$ symmetry. Indeed, if we set $z_{\bm{x}}^{a}=\sigma^{a}_{\bm{x}}+i\sigma^{a+N}_{\bm{x}}$, where $\sigma^{a}_{\bm{x}}$ is a real $2N$-component vector, we obtain
$H=-JN\sum_{{\bm{x}},\mu}\sigma_{\bm{x}}\cdot\sigma_{{\bm{x}}+\hat{\mu}}.$ (123)
@@
Taking into account the fact that for $\kappa=0$ and $m^{2}=0$ the LAH model reduces itself to a lattice CPN-1 model, see Sec. 3.2.1, and therefore there is a corresponding finite-temperature transition for $N\geq 2$, and no transition for $N=1$, one arrives at the phase diagrams reported in the two upper panels of Fig. 25. For $N=1$ [phase diagram (a)] there is a single transition line, while for $N\geq 2$ [phase diagram (b)], there are three different transition lines intersecting in a single point. For small values of $m^{2}$ the GSB perturbation is not strong enough to induce long-range order in the vector correlations of ${\bm{z}}_{\bm{x}}$, while the gauge invariant spin-2 tensor operator in Eq. (60) is ordered, as for $m^{2}=0$. We expect the CPN-1 behavior to extend along the D-T transition line, which starts at $m^{2}=0$, as shown in Ref. [501] for a general GSB perturbation of the compact model, and verified numerically in Ref. [302]. In particular, for $N=2$, continuous transitions should belong to the O(3) universality class. The irrelevance of the GSB perturbation for small $m^{2}$ is not unexpected, since gauge modes do not play any active role at the CPN-1 transition. For large $m^{2}$ values, in the low-temperature phase also vector correlations become long ranged. A natural hypothesis is that the critical behavior observed along the T-TV and D-TV transition lines, for finite values of $J$ and $m^{2}$, is the same as that observed for $J\to\infty$ and $m^{2}\to\infty$, respectively. Thus, continuous transitions should belong to the XY and O($2N$) universality classes, respectively. The phase diagrams reported in the two upper panels of Fig. 25 are supposed to hold also for small values of $\kappa$. On the other hand, for large values of $\kappa$, we expect a different behavior.
@@
Before discussing the large-$\kappa$ behavior, let us consider the limit $J\to\infty$ for generic $\kappa$ and $m^{2}$. Since in this limit the gauge fields take the form (124), the effective Hamiltonian is equivalent to an IXY model (discussed in Sec. 3.5.1) coupled with the Villain XY model (125). In panel (c) of Fig. 25 we show the expected phase diagram, showing a transition line connecting the Villain XY model transition for $\kappa=0$ with the transition of the IXY model on the $m^{2}=0$ axis. Continuous transitions should belong to the XY vector universality class for $m^{2}\not=0$. While there is an XY transition if $m^{2}$ is varied for fixed small $\kappa$, no transition occurs when $\kappa$ is fixed to a large value. Thus, for large $\kappa$ we expect a single transition line, see panel (d) of Fig. 25, connecting the $m^{2}=0$ transition to the O($2N$) transition occurring for $m^{2}\to\infty$. It is natural to expect that continuous transitions along the whole line belong to the O($2N$) universality class for $m^{2}>0$. In particular, the charged transitions that occur for $m^{2}=0$ and $N>N^{*}$ should be unstable with respect to the addition of the GSB term.
@@
We finally consider the expected phase diagrams for fixed small values of $m^{2}$, see panels (e) and (f) of Fig. 25. They are qualitatively similar to the phase diagrams occurring for $m^{2}=0$, see Fig. 2. However, the nature of the phases, and therefore of the transitions, is different. The Coulomb phase is replaced by a disordered phase, in which both scalar and gauge modes are gapped, while in the low-temperature phase both scalar and gauge fields are ordered. Note that, in the latter phase, because of the absence of gauge invariance, both the gauge-invariant tensor correlations and the non-gauge-invariant vector correlations are long ranged for $N\geq 2$. In the molecular phase (which is present only for $N\geq 2$), the addition of the GSB term changes the nature of the gauge excitations—the Coulomb phase is replaced by a gapped gauge-field phase—but not that of the scalar modes: tensor correlations display long range order and vector correlations are disordered, as in the gauge-invariant model. As for the transition lines, the small-$\kappa$ transition line (which is only present for $N\geq 2$) corresponds to CPN-1 transitions as for $m^{2}=0$: the GSB term is an irrelevant perturbation. On the other hand, the nature of the other two transitions changes. The topological IXY transitions are now replaced by standard XY vector transitions, while the continuous charged transitions that occur for large values of $N$ become O($2N$) vector transitions.
@@ 11.2 Gauge-symmetry breaking in lattice Abelian Higgs models with compact gauge fields
The effect of adding a GSB perturbation to the unit-charge ($Q=1$) LAH model with compact gauge variables, see Sec. 6.1, is discussed in Refs. [302, 501]. We consider a GSB term that is analogous to the photon-mass term in the small gauge limit, i.e.,
$H_{m}=-w\sum_{{\bm{x}},\mu}\hbox{Re }\lambda_{{\bm{x}},\mu}.$ (126)
@@
The behavior of this combined model is analogous to the one presented in the noncompact case. Indeed, for $w\to\infty$, we can set $\lambda_{{\bm{x}}\mu}=1$, and thus we obtain again the O($2N$) vector model (123). On the other hand, for $J\to\infty$, $\lambda_{{\bm{x}},\mu}=1$ modulo gauge transformations, i.e., $\lambda_{{\bm{x}},\mu}=e^{i(\phi_{\bm{x}}-\phi_{{\bm{x}}+\hat{\mu}})}$. Substituting this expression in the GSB term (126), we obtain the effective Hamiltonian
$H_{m}=-w\sum_{{\bm{x}},\mu}\cos(\phi_{\bm{x}}-\phi_{{\bm{x}}+\hat{\mu}}),$ (128)
@@
We now present an argument—it rephrases a similar argument of Ref. [41]—that shows that the phase behavior discussed above generally holds for generic GSB perturbations at transitions where gauge fields are not critical. Indeed, let us consider the partition function of a generic model with a GSB perturbation that only depends on the gauge fields:
$Z=\int[dU_{{\bm{x}},\mu}]\,[d\Phi_{\bm{x}}]\,\exp[-\beta H(U_{{\bm{x}},\mu},\Phi_{\bm{x}})],\qquad H(U_{{\bm{x}},\mu},\Phi_{\bm{x}})=H_{\rm GI}(U_{{\bm{x}},\mu},\Phi_{\bm{x}})+H_{\rm GSB}(U_{{\bm{x}},\mu}),$ (129)
where $H_{\rm GI}$ is the gauge-invariant Hamiltonian, and $H_{\rm GSB}$ is a generic Hamiltonian term breaking the gauge symmetry, depending only on the link gauge variables $U_{{\bm{x}},\mu}$. We now perform a change of variables on the scalar and gauge fields that corresponds to a gauge transformation—thus $Z$, or any expectation value of gauge-invariant operators, does not change. In particular, we redefine $U_{{\bm{x}},\mu}\to V_{{\bm{x}}}U_{{\bm{x}},\mu}V^{\dagger}_{{\bm{x}}+\mu}$. As $H_{\rm GI}$ is gauge invariant, the partition function becomes
$Z=\int[dU_{{\bm{x}},\mu}][d\Phi_{\bm{x}}]\,\exp\left[-\beta H_{\rm GI}(U_{{\bm{x}},\mu},\Phi_{\bm{x}})-\beta H_{\rm GSB}(V_{{\bm{x}}}U_{{\bm{x}},\mu}V^{\dagger}_{{\bm{x}}+\mu})\right].$ (130)
@@
The new Hamiltonian $\widehat{H}$ is clearly gauge invariant, and equivalent to the original one, if we consider the partition function and, more generally, any gauge-invariant correlation function. The Hamiltonian $\widehat{H}$ contains interactions between the gauge variables $U_{{\bm{x}},\mu}$ and $U_{{\bm{y}},\nu}$ at any distance $|{\bm{x}}-{\bm{y}}|$. However, if these interactions are exponentially suppressed for $|{\bm{x}}-{\bm{y}}|\to\infty$, then $\widehat{H}$ represents a gauge-invariant model with effective short-range interactions. This is the general scenario that we expect to emerge when the gauge fields are not critical. This can be verified by performing a strong-coupling expansion assuming $\beta w$ small. For example, in the case of the GSB perturbation (126), $\widehat{H}$ can be written as a sum of lattice loops. In the strong-coupling expansion, a lattice loop of length $L$ is weighted by a factor that behaves as $(\beta w)^{L}$ for $\beta w\to 0$. For instance, the leading term is the plaquette, with a weight of order $(\beta w)^{4}$, which renormalizes the value of $\gamma$. The next term corresponds to the $2\times 1$ plaquette, with a coefficient proportional to $(\beta w)^{6}$, and so on. Couplings therefore scale as $\exp(-a|{\bm{x}}-{\bm{y}}|)$ with $a\sim-\ln(\beta w)$, showing the short-range nature of the interactions.
This argument proves that, for sufficiently small, but finite, values of the GSB parameter $w$, the partition function and the gauge-invariant correlation functions can be computed in an equivalent gauge-invariant theory, without GSB terms, with short-range interactions. Finally, to conclude the argument, let us note that we are considering a model in which gauge fields do not play any role, i.e., the critical behavior is independent of the gauge-field interactions, and therefore it is the same as in the original model with $w=0$. We conclude that the phase structure is independent of $w$ for sufficiently small values. Note that the argument does not rely on the Abelian nature of the theory and thus is should also hold in non-Abelian gauge models.
We finally mention that the same arguments also apply to a different model without gauge fields. Indeed, for $\kappa=0$, one can integrate out the link variables $\lambda$, obtaining an effective Hamiltonian
@@
with $\widehat{w}=w/(2JN)$. As it occurs in the CPN-1 model, we expect the Hamiltonian $H_{{\rm eff}}$ to have the same critical behavior as the Hamiltonian obtained by replacing the Bessel function $I_{0}(x)$ with its small-$x$ expansion. Since